|
ID |
Date |
Author |
Subject |
|
292
|
Sun Feb 27 01:46:56 2022 |
David, Pierre-Michell | Current status | So far looks like with a temperature of 95K the target is maintained with a reasonable density throughout long periods of time.
Ring current 6.41e7 particles/s
Merger, Tape server (WS3): 1472 kB/s writing fine
Leakage current (WS4): 3.26 uA
DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed
Online monitoring (WS 5): everything fine |
|
291
|
Sun Feb 27 00:53:24 2022 |
David, Pierre-Michell | Current status | We have recovered the beam at 01:49, after such a long time we had no target so we increased from 80K to 120K at 01:52. 02:01 update 120K-->95K as we have observed around that temperature always fluctuations and is safer.
Ring current 7.74e7 particles/s
Merger, Tape server (WS3): 1619 kB/s writing ok
Leakage current (WS4): 3.26 uA
DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed
Online monitoring (WS 5): done
|
|
290
|
Sun Feb 27 00:07:15 2022 |
David, Pierre-Michell | No beam, valve closed | At 00:37 we lost the beam due to a high pressure burst in the tarjet which resulted in a valve closed at YR09VV...
On-call service technicians are coming to solve the problem.
|
|
289
|
Sat Feb 26 23:20:13 2022 |
David, Pierre-Michell | Update status | In this last hour we have been playing with the cryostat temperature as dictated by Nikos, in order to have the target working properly. For this, we have increased from 80K to 140K the cryostat temperature then we saw again target until we didn't. Then we went down again to 80K. After a few minutes, with a really nice target again was gone, so we increased the temperature to 120K at 00:19. At 00:27 decrease from 120K to 80K.
Ring current 8e7 particles/s
Merger, Tape server (WS3): 1580 kB/s writing ok
Leakage current (WS4): 3.27 uA
DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed
Online monitoring (WS 5): done |
|
288
|
Sat Feb 26 22:22:27 2022 |
David, Pierre-Michell | Starting to take data | At 23:22 we started recording data under the new conditions.
Merger, Tape server (WS3): 1408 kB/s writing ok
Leakage current (WS4): 3.27 uA
DAQ (WS2): temperature checked, clock status:ok, adc calibration:ok, white rabbit check:failed
Online monitoring (WS 5): done |
|
287
|
Sat Feb 26 04:36:38 2022 |
RSS TD | Saturday 26 February 07:00-15:00 | 05:36 E4 pressure c. 1.7e-9mbar *not* 3e-10mbar as reported earlier
See attachments 1 & 2
E4 pressure not < 5e-10mbar - wait to startup gas jet target
06:47 RSS & TD on-site
07.00 E4 1.70e-9mbar
S3 9.42e-10mbar
07:00 Checks: all good except for the white rabbit decoder status.
07:15 All histograms cleared in the run control tab.
07.32 E4 1.60e-9mbar
S3 9.00e-10mbar
08:06 E4 1.49e-9mbar
S3 8.59e-10mbar
08:30 E4 1.43e-9mbar
S3 8.34e-10mbar
09:00 E4 1.35e-9mbar
S3 8.03e-10mbar
09:30 E4 1.29e-9mbar
S3 7.70e-10mbar
10:00 E4 1.23e-9mbar
S3 7.49e-10mbar
10:00 Checks: all good except for the white rabbit decoder status.
10:30 E4 1.18e-9mbar
S3 7.22e-10mbar
11:00 E4 1.14e-9mbar
S3 7.02e-10mbar
11:20 Compressor was switched on.
11:30 E4 1.10e-9mbar
S3 6.82e-10mbar
12:00 E4 1.35e-9mbar
S3 6.64e-10mbar
12:30 E4 1.23e-9mbar
S3 6.46e-10mbar
14:45 Checks: all good except for the white rabbit decoder status.
17.20 System wide checks OK *except* WR decoder status
Base Current Difference
aida01 fault 0xab60 : 0xab85 : 37
aida02 fault 0x88d3 : 0x88f7 : 36
aida03 fault 0x23aa : 0x23cf : 37
aida04 fault 0x15d7 : 0x15fc : 37
aida05 fault 0x2f03 : 0x2f0e : 11
White Rabbit error counter test result: Passed 0, Failed 5
FEE64 temperatures OK - attachment 26
stats (adc data items) OK - attachment 27
DSSSD bias -100V leakage current -3.31uA |
|
286
|
Fri Feb 25 19:20:22 2022 |
MS, PJW | check | 'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status' no changes
|
|
285
|
Fri Feb 25 17:48:22 2022 |
MS, PJW | check | 'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status' - no changes
|
|
284
|
Fri Feb 25 16:01:04 2022 |
MS, PJW | check |
'Check Clock Status': Clock status test result: Passed 5, Failed 0
'Check ADC Calibration' : Calibration test result: Passed 5, Failed 0
'Check the White rabbit decoder status':
Base Current Difference
aida01 fault 0xab60 : 0xab7f : 31
aida02 fault 0x88d3 : 0x88f2 : 31
aida03 fault 0x23aa : 0x23c9 : 31
aida04 fault 0x15d7 : 0x15f6 : 31
aida05 fault 0x2f03 : 0x2f0c : 9
White Rabbit error counter test result: Passed 0, Failed 5
histograms was reset. |
|
283
|
Fri Feb 25 15:38:40 2022 |
JM, CB, PJW | Run 55, 56, 59 root histograms | Run55 - detector out, start 22/2/22 13:20, end 23/2/22 10:49
Run 56 - detector moved in, start 23/2/22 10:4, end 23/2/22 15:46
Run 59 - detector position optimised, detector position in, start 23/2/22 17:57, end 24/2/22 00:04 |
|
282
|
Fri Feb 25 11:20:30 2022 |
CB | Pressure and temperature data saved | Saved pressure reading and temperature data logs from internal target in /TapeServer/22Feb/OtherData
Leakage current data is saved in /home/npg/Programs/caenlogger
|
|
281
|
Fri Feb 25 09:58:08 2022 |
GHC, CB, ML, EM | Current status | 10:58
Started new run: 77
Nozzle clogged more when trying warming again.
Warming up, venting top and bottom to clean nozzle.
|
|
280
|
Fri Feb 25 08:14:47 2022 |
George, Esther | Current status | 9:14
Beam intensity around ~7.5e7
Pressures OK
System check OK
White rabbit fail diff: 24
Write speed: ~1.8Mb/s
Leakage current 3.49uA
Cryostat Setpoint: 121.5K |
|
279
|
Fri Feb 25 06:57:09 2022 |
George, Esther | Current status | 08:00
Beam seems to be back
Intensity ~7.8e7
Pressures OK
System check OK
White rabbit fail diff: 24
Write speed: ~1.3Mb/s
Leakage current 3.49uA
Cryostat Setpoint: 121.5K |
|
278
|
Fri Feb 25 06:28:12 2022 |
George, Esther | Current Status | 07:30
Some signals on the beam monitor, seems they are optimising it.
System Check OK
Temps OK
White Rabbit fail diff: 24
Leakage Current: 3.48uA
Pressures OK
Data rate ~1Mb/s |
|
277
|
Fri Feb 25 05:04:16 2022 |
Rui-Jiu Chen, Alex | check | Still no beam in CRING.
Motor fully out position.
SW3
- Merger OK
- Screenshot uploaded
- Data rate : 217 kbytes/s
SW4
- current plot uploaded
SW2
- T OK
- system wide checks OK
- Check statistics tab ok
- Spectra ok
SW5
4 plots uploaded |
|
276
|
Fri Feb 25 03:20:29 2022 |
Rui-Jiu Chen, Alex | Current status | Gleb Vorobyev tried for few hours to fix the ion source but without success.
So, there will be no beam in CRYRING during this night shift. Today at about 08:00 the efforts to fix the ion source will be restarted. |
|
275
|
Fri Feb 25 03:01:11 2022 |
Rui-Jiu Chen, Alex | check | Still no beam in CRING.
Motor fully out position.
SW3
- Merger OK
- Screenshot uploaded
- Data rate : 236 kbytes/s
SW4
- current plot uploaded
SW2
- T OK
- system wide checks OK
- Check statistics tab ok
- Spectra ok
SW5
4 plots uploaded |
|
274
|
Fri Feb 25 01:12:49 2022 |
Rui-Jiu Chen, Alex | Check | Still no beam in CRING.
Motor fully out position.
SW3
- Merger OK
- Screenshot uploaded
- Data rate : 243 kbytes/s
SW4
- current plot uploaded
SW2
- T OK
- system wide checks OK
- Check statistics tab ok
- Spectra ok
SW5
4 plots uploaded |
|
273
|
Fri Feb 25 00:00:40 2022 |
Rui-Jiu Chen, Alex | Check | There are some problems with ion source. No beam in CRYING.
Motor fully out position.
SW3
- Merger OK
- Screenshot uploaded
- Data rate : 230 kbytes/s
SW4
- current plot uploaded
SW2
- T OK
- system wide checks OK
- Check statistics tab ok
- Spectra ok
SW5
4 plots uploaded |
|